OpenStudy (hellokitty17):

MEDAL FOR RIGHT ANSWER HERE!!!!!!!!!!!!!!!! A biologist created the following graph to show the relationship between the temperature of water (x), in degrees Celsius, and the number of insect larvae (y) in the water: graph of y equals minus 2 times the square of x plus 20 times x plus 400 What do the x-intercepts of the graph represent? There were 20 larvae at 0 degrees Celsius. There were 10 larvae at 0 degrees Celsius. The water has no larvae at -10 degrees Celsius and 20 degrees Celsius. The water has maximum larvae at -10 degrees Celsius and 20 degrees Celsius.

OpenStudy (hellokitty17):

@Mehek14

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} y=-2x^{2}+20x+400\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (hellokitty17):

what does that mean

OpenStudy (mathmath333):

for \(x\)-intercepts set \(y=0\) and solve \(\large \color{black}{\begin{align} 0=-2x^{2}+20x+400\hspace{.33em}\\~\\ \end{align}}\)

oregonduck (oregonduck):

and the answer is?

OpenStudy (mathmath333):

do u know how to solve this \(\large \color{black}{\begin{align} 0=-2x^{2}+20x+400\hspace{.33em}\\~\\ \end{align}}\)

oregonduck (oregonduck):

no

oregonduck (oregonduck):

A biologist created the following graph to show the relationship between the temperature of water (x), in degrees Celsius, and the number of insect larvae (y) in the water: graph of y equals minus 2 times the square of x plus 20 times x plus 400 ..........y = -2x² + 20x + 400 What do the x-intercepts of the graph represent? ..........C) The water has no larvae at -10 degrees Celsius ................and 20 degrees Celsius.